A uniform beam 6m long and weighing 4kg rests on supports P and Q placed left and right 1M from each end of the beam. Weights of mass 10kg and 8kg are placed near P and Q respectively, right at each end of the beam. Calculate the reactions at P and Q.


1
Expert's answer
2021-03-23T08:08:09-0400

Take moments about P:


"(10)(9.8)(1)+Q(4)=(4)(9.8)(2)+(8)(9.8)(5)\\Q=93.1\ N"

Take moments about Q:


"(10)(9.8)(5)+(4)(9.8)(2)=4P+(8)(9.8)(1)\\P=122.5\ N"